For a two-tailed test with variance unknown, n=19, and a=0.05, what is the critical value?

1
Expert's answer
2022-05-31T07:33:06-0400

This corresponds to a two-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha = 0.05," "df=n-1=18" and the critical value for a two-tailed test is "t_c = 2.100922."